40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 24 03:58:13 2012, kiwamu, Update, LSC, Y arm + PRMI part II 
    Reply  Fri Feb 24 15:01:31 2012, kiwamu, Update, LSC, Y arm + PRMI part II time_series.png
       Reply  Fri Feb 24 19:18:28 2012, kiwamu, Update, LSC, Y arm + PRMI : how they should look like expected_time_series.png
    Reply  Fri Feb 24 18:37:13 2012, rana, Update, LSC, Y arm + PRMI part II 
Message ID: 6313     Entry time: Fri Feb 24 15:01:31 2012     In reply to: 6310     Reply to this: 6317
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Y arm + PRMI part II 

The figure below shows the time series of the Y arm + PRMI trail.

time_series.png

(Top plot )

  Normalized TRY (intracavity power). It is normalized such that it shows 1 when the arm is locked with the recycling mirrors misaligned.

(Middle plot)

ASDC and REFLDC in arbitrary unit.

(Bottom plot)

The amount of the arm length detuning observed at the fine frequency discriminator.

 

(Sequence)

At t = 20 sec, the amount of detuning was adjusted so that the cavity power goes to the maximum. At this point the PRM was misaligned.

At t = 30 sec, the cavity length started being slowly detuned to 10 nm. As it is being detuned the intracavity power goes down to almost zero.

At t = 45 sec, the alignment of PRM was restored. Because of that, the REFLDC and ASDC diodes started receiving a large amount of light.

At t = 85 sec, the PRCL and MICH were locked. The REFLDC signal became a high value as the carrier light is mostly reflected. The ASDC goes to a low value as the MICH is kept in the dark condition.

At t = 100 sec, the length started being slowly back to the resonance while the PRMI lock was maintained.

At t = 150 sec, the lock of the PRCL and MICH were destroyed. With the arm fully resonance, I wasn't able to recover the PRMI lock with the same demod signals.

Quote from #6310

I tried the Yarm + PRMI configuration again.

The PRMI part was locked, but it didn't stay locked during the Y arm was brought to the resonance point.

I will post the time series data later.

 

ELOG V3.1.3-